# A compact model of Escherichia coli core and biosynthetic metabolism

### v1

@@ 2.5 Enzyme-constrained Flux Balance Analysis with EC-iCH360
To increase the predictive ability of the model, we thus adjusted the turnover numbers by fitting them to experimental measurements through a custom heuristic (Methods, Supplementary Information A.5). By simultaneously fitting all available conditions, we ensured that our adjustment procedure is robust to condition-specific biases. Further, by introducing regularisation within our adjustment scheme, we penalised large deviations of parameters from the original dataset, increasing the robustness of the procedure to overfitting. Our adjusted parameter set shows a mean absolute deviation (computed for log10-transformed turnover values) from the original parameter set of $\approx 0.22$ (Supplementary Figure 10) and results in significantly better agreement with experimental measurements across conditions (Figure 4A). Further, mean enzyme abundances across conditions are very well predicted with the adjusted parameter set (Figure 4B), implying that residuals between measurements and predictions are now to be attributed to variability in enzyme cost across conditions (which simple frameworks such as enzyme-constrained FBA cannot account for), rather than systematic over- or under-predictions. Thus, each adjusted turnover parameter can be thought of as a ”typical” apparent $k_{\rm cat}$ value, incorporating average saturation trends for an enzyme across growth conditions.
@@ 4.7 Adjustment of turnover numbers across conditions
Turnover numbers from [heckmann_kinetic_2020] were adjusted based on the nonlinear programming (NLP) formulation detailed in Supplementary Information A.5). The reference flux distributions required by the procedure were obtained using the original (unadjusted) set of turnover numbers and the bounds on allowable adjustments ($\mathbf{l}$ and $\mathbf{u}$ in Eq. 29) set to ±$\pm 2$ (corresponding to a maximal 100-fold increase or reduction of each parameter from the original value). The linear program used to obtain reference flux distributions was formulated and solved with GUROBI [gurobi], while the nonlinear program used for turnover adjustment was formulated and solved with the open-source optimisation package CasADi [Andersson2019]. To investigate the effect of the ridge regularisation hyperparameter ($\rho$ in equation 29), we solved the adjustment problem for a broad range of values of this parameter and, each time, we computed the RMSE between measurements and predictions of enzyme abundance, as well as the mean absolute deviation of between original and adjusted turnover values, both computed for log10-transformed data (Supplementary Figure 11). Based on this information, a value of $\rho=1$, after which any further decrease in the amount of regularisation results in marginal reduction of the RMSE, was chosen to compute the final set of adjusted turnover numbers. Upon reparametrisation into apparent turnover numbers, (see Supplementary Information A.5.3), this set of adjusted parameters was used to parametrise the enzyme-constrained model variant EC-iCH360.
@@ 4.8 Enzyme allocation predictions
To validate enzyme allocation predictions against experimental values, we first retrieved measured polypeptide abundances for each growth condition from [schmidt_quantitative_2016] and imputed missing values using, whenever available, abundance values from the PAX Database [huang_paxdb_2023]. Then we estimated enzyme counts across conditions from polypeptide counts using non-negative least-squares estimation (Supplementary Information A.4). Next, we converted them into mass abundances (in units of ${\rm g~{}gDW^{-1}}$) based on the molecular mass of each complex (see Supplementary Information A.3) and assuming a cell mass of $2.8\times 10^{-13}\rm{g}$ (BIONUMBER ID 103904 [milo_bionumbers--database_2010]). Enzyme allocation predictions for each condition were then computed via EC-iCH360 by fixing the growth rate to the experimentally measured one from [schmidt_quantitative_2016] and minimising the total enzyme cost. To compute predictions using the turnover number estimates from [heckmann_kinetic_2020], the average saturation coefficient $\sigma$ was estimated, for each condition as:
@@ 4.12 Component contribution estimates of thermodynamic constants
Estimates of the free energies of reactions, and their uncertainties, were obtained using the component contribution framework previously described in [noor_consistent_2013]. Several reactions in the model involve protein side groups, such as the acyl-carrying protein (ACP), or cofactors, such as glutaredoxin, for which a decomposition in terms of chemical groups is not available. As a result, thermodynamic constants for these reactions cannot be directly estimated through database-based implementations of the component contribution method, such as eQuilibrator [beber_equilibrator_2022]. However, since these non-decomposable protein groups are conserved in all reactions within our model, their net contribution to the reaction thermodynamics is, at least from a group contribution perspective, null. Hence, we first augmented the group incidence matrix of the eQuilibrator database (see Section S1.1 in [beber_equilibrator_2022]) by treating protein groups as non-decomposable units and manually curated the group decomposition of the metabolites containing them. The mean and covariance of the Gibbs standard free energy of reaction ($\Delta_{r}G^{\circ}$) estimates for all reactions in the model were then computed using this augmented matrix as described in [beber_equilibrator_2022].
In order to compute transformed Gibbs free energies of reactions ($\Delta_{r}G^{\prime\circ}$), for which an exact chemical definition of each metabolite is required, protein side groups were replaced by an appropriate chemical moiety best approximating the chemical environment of the metabolite. Specifically, ACP- groups were replaced by a phosphopantetheine group, the natural prosthetic group of acyl-carrier proteins, with a methyl group at the attachment site to the protein scaffold. Similarly, glutaredoxin was replaced by its Cys-Pro-Tyr-Cys active site, with the two cysteines being either free (for the reduced form of the cofactor) or linked by a disulfide bridge (for the oxidised state).
@@ 6 Competing interests
The authors declare no competing interest.
@@ 7 Code availability
The model, together with all relevant data, is available on Github at https://github.com/marco-corrao/iCH360. The code and files required to reproduce all analysis in this manuscript are available on Github at https://github.com/marco-corrao/iCH360_paper.
@@ A.5 Adjustment of turnover numbers based on proteomics measurements across conditions
In this section, we outline the procedure used to adjust the turnover numbers in EC-iCH360 by fitting proteomic measurements across conditions (see Section 2.5 in the main text). Briefly, our aim is to adjust the turnover numbers that parametrise the model so that enzyme allocation predictions obtained through the enzyme-constrained formulation of FBA (see Methods in the main text) more closely match experimental measurements of enzyme abundances. By simultaneously fitting experimental measurements across many growth conditions, we improve the robustness of the fitting procedure to experimental error and generate a condition-independent parameter set that predicts average trends of enzyme allocation across conditions.
@@
Adjusting turnover numbers is, in general, not simple. Due to the linear programming formulation underlying the enzyme-constrained FBA problem, optimal flux distributions (and, by direct consequence, enzyme allocation predictions) are discontinuous over the turnover parameter space, making derivative-based searches through this space problematic from a numerical perspective. Similarly, the high dimensionality of the parameter space limits the applicability of gradient-free optimisation algorithms. Hence, we restrict ourselves to the (comparatively simpler) problem of adjusting turnover parameters for a fixed set of reference flux distributions across growth conditions, constraining our search to the portion of parameter space in which these reference flux vectors remain optimal for their respective conditions. Our adjustment heuristic thus consists of two steps, which we detail in the remainder of this section
A.5.1 Obtaining a set of reference flux distributions
@@
$\begin{cases}\frac{M_{i}}{\sigma_{j}~{}k_{\text{cat},i}}\quad&\text{if reaction $i$ is enzymatic }\\ 0\quad&\text{otherwise}\end{cases}$ (21)
Here, $M_{i}$ is the molecular weight of the enzyme associated with the reaction, $k_{\text{cat},i}$ is the turnover number for the enzyme-reaction pair, and $\sigma_{j}$ is a condition-specific coefficient (typically interpreted as an average enzyme saturation level) scaling the total enzyme abundance predicted for a flux distribution. Given an initial vector of turnover numbers to be adjusted, we compute a reference flux distribution for the $j$th growth condition, $\mathbf{v}_{j}^{*}$ by fixing the biomass flux, $v_{\rm BM}$ to the experimentally measured rate and minimizing the total enzyme cost:
@@
Here, the objective $\mathbf{a}_{j}^{\top}\mathbf{v}$ is the total enzyme cost for this condition, $\bar{v}_{\mathrm{BM},j}$ is the experimentally measured growth rate for the $j$th growth condition, ${\rm\mathbf{S}}\in\mathbb{R}^{\rm{M}\times\rm{N}}$ is the stoichiometric matrix of the network and ${\rm\mathbf{B}}_{j}\in\mathbb{R}^{\rm{P}\times\rm{N}}$ and $\mathbf{b}_{j}\in\mathbb{R}^{\rm{P}}$ are a matrix and a vector, respectively, encoding any desired upper bound (or positive lower bound) on the fluxes for the $j$th growth condition. Noting that constraint 22c can be equivalently cast as a double inequality, we rewrite the problem in the more general form:
@@
where the biomass flux is assumed to be the last component of the flux vector and the augmented matrices:
@@
were introduced. Finally, from (21) we note that the solution of problem (23) is independent of the choice of the coefficient $\sigma$, as this merely amounts to a scaling of the objective function. Hence, without loss of generality, we can simply set $\sigma=1$ for all growth conditions in this step and fit a condition-specific value for this parameter in the next half of the procedure. Solving (23 for each of the $J$ growth conditions available in the experimental dataset, we obtain a set of optimal flux distributions $\mathcal{V^{*}}=\{\mathbf{v}_{1}^{*},\cdots,\mathbf{v}_{J}^{*}\}$, which we will use as a reference in the next step.
A.5.2 Adjusting turnover numbers for the reference set of flux distributions
@@
We now turn to fit the turnover and saturation parameters parameters against experimental measurements of enzyme abundances. For this purpose, we denote with $\mathbf{p}$ a vector of log10 turnover parameters (one for each enzyme-reaction pair in the model) and with $\mathbf{s}$ a vector of condition-specific log10-scaling factors (i.e. $p_{i}=\log_{10}({k_{\rm{cat},i}})$ and $s_{j}=\log_{10}({\sigma_{j}})$). Further, we denote with $\bar{\mathbf{p}}$ the vector of original log-turnover numbers (i.e. the one used to obtain the reference flux distributions). For the choice of reference flux distribution computed in the previous step, the abundance of the $i$th enzyme in condition $j$, $e_{ij}$, is then a function of $\mathbf{p}$ and $\mathbf{s}$:
$e_{ij}(\mathbf{p},\mathbf{s})=\sum_{k}a_{kj}v^{*}_{kj}$ (25)
@@
where the summation index $k$ runs across all reactions catalysed by enzyme $i$ and the flux cost of reaction $k$ in condition $j$, $a_{kj}$, is computed as in (21). From the formulation of the linear program (23), it is clear that there must exist a region $\mathcal{S}_{j}$ of log-turnover parameter space such that $\bar{\mathbf{p}}\in\mathcal{S}_{j}$ and that, for every $\mathbf{p}\in\mathcal{S}_{j}$, $\mathbf{v}_{j}^{*}\in\mathcal{V}^{*}$ is the optimal solution of problem (23) for its growth condition. Hence, in this step of the adjustment procedure, we aim to find a set of adjusted log turnover parameters, $\mathbf{p}^{*}$ and log scaling factors, $\mathbf{s}^{*}$ by minimising the discrepancy between enzyme abundance predictions and measurements, constraining our search to this region of the parameter space $\mathcal{S}\equiv\bigcap\limits_{j}\mathcal{S}_{j}$ where all reference flux distributions are optimal for their respective growth condition:
@@
The objective function of the nonlinear program (26) is a combination of two terms. The first term penalises the mean squared deviation between measurements and predictions of log-enzyme abundance. Note that the above definition of $l(\cdot)$ implies that, for each condition, only enzymes with nonzero predicted abundance are included in this term. The second term is a regularisation expression (whose strength is controlled by the hyperparameter $\rho$) penalising mean squared deviation from the original parameter set. The latter term mainly serves two purposes: first, it ensures that, whenever a turnover parameter is ”free” in the problem (which can happen if its associated reaction fluxes are 00 across all conditions, or if no experimental measurements are available for its associated enzyme), it will be kept at its original value; secondly, it provides a mean to tune the strength of the adjustment procedure.
In order to define the region $\mathcal{S}$, we shall exploit the sufficient optimality conditions of LP (23). Introducing, for each growth condition, the vectors of dual variables, $\boldsymbol{\lambda}_{j}\in\mathbb{R}^{M}$ and $\boldsymbol{\mu}_{j}\in\mathbb{R}^{P}$, corresponding to constraints 23b and 23c, respectively, we note that problem 23 admits (for the $j$th growth condition) a dual problem in the form:
@@
A well-known result in linear programming duality theory [maranas_zomorrodi_2016] states that a flux distribution is the optimal solution of the primal problem 23, if, and only if, the dual problem 28 is feasible (dual feasibility) and its optimal objective coincides with the primal optimal objective, that is $\mathbf{a}_{j}^{\top}\mathbf{v}_{j}^{*}=-\mathbf{b}_{j}^{\top}\boldsymbol{\mu}_{j}$ (strong duality). Taken together, dual feasibility and strong duality thus define the region of optimality in turnover parameter space of each reference flux distribution. Hence, we can integrate the above definition of $\mathcal{S}$ within problem 26 by introducing the two vectors of dual variables ($\boldsymbol{\lambda}_{j}$ and $\boldsymbol{\mu}_{j}$) for each condition as additional optimisation variables, and simultaneously enforcing the optimality for each reference distribution. By doing this, we obtain the final formulation of the nonlinear program for turnover number adjustment, which we solved to obtain the results shown in the main text:

### v2

@@ 2.5 Enzyme-constrained Flux Balance Analysis with EC-iCH360
To increase the predictive ability of the model, we thus adjusted the turnover numbers by fitting them to experimental measurements through a custom heuristic (Methods, Supplementary Information A.5). By simultaneously fitting all available conditions, we ensured that our adjustment procedure is robust to condition-specific biases. Further, by introducing regularisation within our adjustment scheme, we penalised large deviations of parameters from the original dataset, increasing the robustness of the procedure to overfitting. Our adjusted parameter set shows a mean absolute deviation (computed for log10-transformed turnover values) from the original parameter set of $\approx 0.22$ (Supplementary Figure 10) and results in significantly better agreement with experimental measurements across conditions (Figure 4A). Further, mean enzyme abundances across conditions are very well predicted with the adjusted parameter set (Figure 4B), implying that residuals between measurements and predictions are now to be attributed to variability in enzyme demand across conditions (which simple frameworks such as enzyme-constrained FBA cannot account for), rather than systematic over- or under-predictions. Thus, each adjusted turnover parameter can be thought of as a ”typical” apparent $k_{\rm cat}$ value, incorporating average saturation trends for an enzyme across growth conditions.
@@ 4.7 Adjustment of turnover numbers across conditions
Turnover numbers from [heckmann_kinetic_2020] were adjusted based on the nonlinear programming (NLP) formulation detailed in Supplementary Information A.5). The reference flux distributions required by the procedure were obtained using the original (unadjusted) set of turnover numbers and the bounds on allowable adjustments ($\mathbf{l}$ and $\mathbf{u}$ in Eq. 35) set to ±$\pm 2$ (corresponding to a maximal 100-fold increase or reduction of each parameter from the original value). The linear program used to obtain reference flux distributions was formulated and solved with GUROBI [gurobi], while the nonlinear program used for turnover adjustment was formulated and solved with the open-source optimisation package CasADi [Andersson2019]. To investigate the effect of the ridge regularisation hyperparameter ($\rho$ in equation 35), we solved the adjustment problem for a broad range of values of this parameter and, each time, we computed the RMSE between measurements and predictions of enzyme abundance, as well as the mean absolute deviation of between original and adjusted turnover values, both computed for log10-transformed data (Supplementary Figure 11). Based on this information, a value of $\rho=1$, after which any further decrease in the amount of regularisation results in marginal reduction of the RMSE, was chosen to compute the final set of adjusted turnover numbers. Upon reparametrisation into apparent turnover numbers, (see Supplementary Information A.5.5), this set of adjusted parameters was used to parametrise the enzyme-constrained model variant EC-iCH360.
@@ 4.8 Enzyme allocation predictions
To validate enzyme allocation predictions against experimental values, we first retrieved measured polypeptide abundances for each growth condition from [schmidt_quantitative_2016] and imputed missing values using, whenever available, abundance values from the PAX Database [huang_paxdb_2023]. Then we estimated enzyme counts across conditions from polypeptide counts using non-negative least-squares estimation (Supplementary Information A.4). Next, we converted them into mass abundances (in units of ${\rm g~{}gDW^{-1}}$) based on the molecular mass of each complex (see Supplementary Information A.3) and assuming a cell dry mass of $2.8\times 10^{-13}\rm{g}$ (BIONUMBER ID 103904 [milo_bionumbers--database_2010]). Enzyme allocation predictions for each condition were then computed via EC-iCH360 by fixing the growth rate to the experimentally measured one from [schmidt_quantitative_2016] and minimising the total enzyme cost. To compute predictions using the turnover number estimates from [heckmann_kinetic_2020], the average saturation coefficient $\sigma$ was estimated, for each condition as:
@@ 4.12 Component contribution estimates of thermodynamic constants
Estimates of the free energies of reactions, and their uncertainties, were obtained using the component contribution framework previously described in [noor_consistent_2013]. Several reactions in the model involve protein side groups, such as the acyl-carrying protein (ACP), or cofactors, such as glutaredoxin, for which a decomposition in terms of chemical groups is not available. As a result, thermodynamic constants for these reactions cannot be directly estimated through database-based implementations of the component contribution method, such as eQuilibrator [beber_equilibrator_2022]. However, since these non-decomposable protein groups are conserved in all reactions within our model, their net contribution to the reaction thermodynamics is, at least from a group contribution perspective, null. If we were only interested in computing standard free energies of reaction, $\Delta_{r}G^{\circ}$, we could simply treat protein groups as non-decomposable ”black-box” units and add them to the group incidence matrix of the eQuilibrator database (see Section S1.1 in [beber_equilibrator_2022]), enabling us to construct a group decomposition for compounds that contain them. However, for the computation of transformed standard free energies of reaction, $\Delta_{r}^{\prime}G^{\circ}$, an exact chemical definition of each metabolite is required.
To address this issue, protein groups were replaced by an appropriate chemical moiety that best approximates the metabolite’s chemical environment. Specifically, ACP- groups were replaced by a phosphopantetheine group, the natural prosthetic group of acyl-carrier proteins, with a methyl group at the attachment site to the protein scaffold. Similarly, glutaredoxin was replaced by its Cys-Pro-Tyr-Cys active site, with the two cysteines being either free (for the reduced form of the cofactor) or linked by a disulfide bridge (for the oxidised state). International Chemical Identifiers (InChI) were constructed for these ”replacement” metabolites and used to extend the default compound cache of the EQuilibrator database. The chemical identifiers for these compounds and the custom-extended EQuilibrator compound cache are available in the repository supporting this manuscript.
@@ 6 Competing interests
The authors declare that they comply with the PCI rule of having no financial conflicts of interest in relation to the content of the article.
@@ 7 Code availability
The model, together with all relevant data, is available on Github at https://github.com/marco-corrao/iCH360. The code and files required to reproduce all analysis in this manuscript are available on Github at https://github.com/marco-corrao/iCH360_paper and on Zenodo at https://doi.org/10.5281/zenodo.11092781.
@@ A.5 Adjustment of turnover numbers based on proteomics measurements across conditions
In this section, we outline the procedure used to adjust the turnover numbers in EC-iCH360 by fitting proteomic measurements across conditions (see Section 2.5 in the main text). Briefly, our aim is to adjust the turnover numbers that parametrise the model so that enzyme allocation predictions obtained through the enzyme-constrained formulation of FBA (see Methods in the main text) more closely match experimental measurements of enzyme abundances. By simultaneously fitting experimental measurements across many growth conditions, we improve the robustness of the fitting procedure to experimental error and generate a condition-independent set of ”typical” apparent turnover numbers that predict average trends of enzyme allocation across conditions. The output of our procedure is a set of typical enzyme efficiencies, one for each enzyme, estimated from proteomic data across conditions, as well as a set of condition-specific scaling factors that account for differences in total measured enzyme abundances across conditions. In section A.5.1 we rigorously define these parameters and state the main assumption underlying our heuristic. In sections A.5.3, A.5.4, and A.5.5 we formulate a two-steps optimisation problem whose solution, upon a suitable reparameterisation, yields data-fitted estimates of the desired parameters.
A.5.1 Preliminaries
Consider enzyme $i$ in a given metabolic state $j$. This enzyme catalyses a metabolic flux $v_{ij}$ given by
$v_{ij}=\kappa_{ij}~{}c^{enz}_{ij}$ (21)
where $c^{enz}_{ij}$ is the enzyme concentration and $\kappa_{ij}$ is the enzyme efficiency (or “apparent turnover number”). The efficiency $\kappa_{ij}$ is a positive rate (here in 1/h, but usually reported in 1/s). Since, by definition, it must be lower than the enzyme’s turnover number $k_{cat,i}$, we write it as
$\kappa_{ij}=\sigma_{ij}~{}k_{cat,i}$ (22)
where $\sigma_{ij}\in[0,1]$ is a unitless “capacity usage” factor. In enzyme-constrained models, enzymes are often expressed by their mass abundance (in g/gDW) instead of enzyme concentrations, so we can write the flux as
$v_{ij}=\frac{1}{a_{ij}}~{}e_{ij}$ (23)
where $a_{ij}$ is the enzyme cost per catalysed flux, given by the enzyme molecular mass $M_{i}$ (in g/mol) divided by the enzyme efficiency $\kappa_{ij}$.
In principle, the capacity usage factors (and therefore efficiencies) of enzymes may freely vary between growth conditions. However, as a heuristics, we here assume that $\sigma_{ij}$ can be approximated as the product of two factors: an enzyme-specific term and a condition-specific one, that is:
$\sigma_{ij}\approx\sigma_{i}\cdot\tau_{j}$ (24)
Here, the enzyme-specific factor $\sigma_{i}$ denotes the ”typical” capacity usage factor of our enzyme in the range of conditions studied. The condition-specific term, $\tau_{j}$, is a unitless scaling factor that simultaneously increases or decreases the efficiencies of all enzymes depending on the cell’s growth conditions. By convention, we assume that the values of $\tau_{j}$ are centered around $1$. Substituting (24) in (22), we obtain an equivalent expression (under our assumptions) for the apparent turnover number $\kappa_{ij}$:
$\kappa_{ij}=\underbrace{\sigma_{i}~{}k_{cat,i}}_{\kappa_{i}}~{}\tau_{j}$ (25)
where $\kappa_{i}$ is a ”typical apparent turnover number” that is condition-independent. Practically, the above assumption allows us to simplify the problem by reducing the number of parameters to be fitted from $I\cdot J$ to $I+J$, where $I$ and $J$ denote the total number of enzymes and conditions considered, respectively. Our heuristic assumption corresponds to the idea that ”high-quality carbon sources” allow a cell to establish metabolic states in which enzyme efficiencies are generally high, allowing for large fluxes per enzyme abundance in all the reactions, and therefore for high cell growth rates. Probably, this heuristics would fail in some other cases, e.g. cases in which enzymes are specifically perturbed by enzyme inhibitors. But in fact, it turns out that our model, assuming a single ”typical apparent turnover number” $k_{app,i}$ for each enzyme, yields very good enzyme allocation predictions. This is exactly what would be expected if our heuristic assumption were correct, and therefore supports our heuristic prediction.
Below we describe how the estimates of (enzyme-specific) efficiencies $\kappa_{i}$ and (condition specific) scaling factors $\tau_{j}$ for our model were obtained from model simulations and proteomics data.
A.5.2 Overview of the fitting procedure
@@
Fitting the typical turnover parameters to proteomic data, in general, not simple. Due to the linear programming formulation underlying the enzyme-constrained FBA problem, optimal flux distributions (and, by direct consequence, enzyme allocation predictions) are discontinuous over the turnover parameter space, making derivative-based searches through this space problematic from a numerical perspective. Similarly, the high dimensionality of the parameter space limits the applicability of gradient-free optimisation algorithms. Hence, we restrict ourselves to the (comparatively simpler) problem of adjusting turnover parameters for a fixed set of reference flux distributions across growth conditions, constraining our search to the portion of parameter space in which these reference flux vectors remain optimal for their respective conditions. This simplified fitting procedure thus consists of two steps. In the first part of the procedure, we use an initial parameter set to compute a set of reference flux distributions (one per growth condition) using enzyme-constrained FBA. In the second part, turnover parameters are fitted based on these reference flux distributions and experimental measurements of enzyme abundances.
A.5.3 Obtaining a set of reference flux distributions
@@
$a_{ij}=\begin{cases}\frac{M_{i}}{\kappa_{ij}}\quad&\text{if reaction $i$ is enzymatic }\\ 0\quad&\text{otherwise}\end{cases}$ (26)
Here, $M_{i}$ is the molecular mass of the enzyme associated with the reaction, $\kappa_{ij}$ is the condition-dependent enzyme efficiency, as defined in 25. Given an initial guess for the value of each $\kappa_{ij}$, we compute a reference flux distribution for the $j$th growth condition, $\mathbf{v}_{j}^{*}$ by fixing the biomass flux, $v_{\rm BM}$ to the experimentally measured rate and minimizing the total enzyme cost:
@@
Here, the objective $\mathbf{a}_{j}^{\top}\mathbf{v}$ is the total enzyme cost for this condition, $\bar{v}_{\mathrm{BM},j}$ is the experimentally measured growth rate for the $j$th growth condition, ${\rm\mathbf{S}}\in\mathbb{R}^{\rm{M}\times\rm{N}}$ is the stoichiometric matrix of the network and ${\rm\mathbf{B}}_{j}\in\mathbb{R}^{\rm{P}\times\rm{N}}$ and $\mathbf{b}_{j}\in\mathbb{R}^{\rm{P}}$ are a matrix and a vector, respectively, encoding any desired upper bound (or positive lower bound) on the fluxes for the $j$th growth condition. Noting that constraint 27c can be equivalently cast as a double inequality, we rewrite the problem in the more general form:
@@
where the biomass flux is assumed to be the last component of the flux vector and the augmented matrices
@@
were introduced.
In order to solve the the linear program 27, we shall consider an initial guess for the turnover parameters and assume that, for each condition, all enzymes operate at the same saturation level, so that $\sigma_{ij}\equiv\bar{\sigma}_{j}$. Since the optimal flux distribution obtained as a solution of problem (28) is unchanged by the choice of $\bar{\sigma}_{j}$ (as this merely amounts to a scaling of the objective function), we can simply set $\bar{\sigma}_{j}=1$ (that is, set $\kappa_{ij}=k_{\rm cat,i}$) for all conditions at this stage.
Solving (28) for each of the $J$ growth conditions available in the experimental dataset, we obtain a set of optimal flux distributions $\mathcal{V^{*}}=\{\mathbf{v}_{1}^{*},\cdots,\mathbf{v}_{J}^{*}\}$, which we will use as a reference in the next step.
A.5.4 Estimating typical enzyme efficiencies for the reference set of flux distributions
We now turn to fitting the relevant parameters against experimental measurements of enzyme abundances. For this purpose, we shall express the the efficiency of each enzyme-condition pair as:
$\kappa_{ij}=\theta_{j}~{}k_{i}$ (30)
where $\theta_{j}$ is a condition-specific scaling term that simultaneously scales all efficiencies for a given condition, while $k_{i}$ is an ”adjusted” turnover parameter that simultaneously accounts for inaccuracies in the original turnover parameter numbers as well as differences in typical saturation across enzymes (and, hence, it’s not formally a turnover number). While the above parameterization differs from the one in equation 25 – the terms have a different interpretation! – it will greatly simplify the notation of the fitting problem, as it allows us to easily distinguish between ”global” adjustments (those affecting all enzymes in a condition), which we wish to pick freely, and ”local” adjustments (affecting the efficiency of individual enzymes), which instead we wish to regularise. While this factorisation of saturation effects is merely a computational convenience and not necessarily biologically meaningful, we will retrieve the parameters in Eq. 25 from those in Eq. 30 upon a simple reparameterisation, as we detail in Section A.5.5
@@ A.5.4 Estimating typical enzyme efficiencies for the reference set of flux distributions
To formulate our fitting procedure, we will denote with $\mathbf{p}$ a vector of log10 adjusted turnover numbers (i.e. $p_{i}=\log_{10}{k_{i}}$) and with $\mathbf{s}$ a vector of condition-specific log10-scaling factor (i.e. $s_{j}=\log_{10}\theta_{j}$). Further, we denote with $\bar{\mathbf{p}}$ the vector of original log-turnover numbers (i.e. the one used to obtain the reference flux distributions). For the choice of reference flux distribution computed in the previous step, the abundance of the $i$th enzyme in condition $j$, $e_{ij}$, is then a function of $\mathbf{p}$ and $\mathbf{s}$:
$e_{ij}(\mathbf{p},\mathbf{s})=\sum_{k}a_{kj}~{}v^{*}_{kj}$ (31)
@@
where the summation index $k$ runs across all reactions catalysed by enzyme $i$ and the flux cost of reaction $k$ in condition $j$, $a_{kj}$, is computed as in (26). From the formulation of the linear program (28), there must exist a region $\mathcal{S}_{j}$ of log-turnover parameter space such that $\bar{\mathbf{p}}\in\mathcal{S}_{j}$ and that, for every $\mathbf{p}\in\mathcal{S}_{j}$, $\mathbf{v}_{j}^{*}\in\mathcal{V}^{*}$ is the optimal solution of problem (28) for its growth condition. Hence, in this step of the adjustment procedure, we aim to find a set of typical log efficiencies, $\mathbf{p}^{*}$ and log scaling factors, $\mathbf{s}^{*}$ by minimising the discrepancy between enzyme abundance predictions and measurements, constraining our search to this region of the parameter space $\mathcal{S}\equiv\bigcap\limits_{j}\mathcal{S}_{j}$ where all reference flux distributions are optimal for their respective growth condition:
@@
The objective function of the nonlinear program (32) is a combination of two terms. The first term penalises the mean squared deviation between measurements and predictions of log-enzyme abundance. Note that the above definition of $l(\cdot)$ implies that, for each condition, only enzymes with nonzero predicted abundance are included in this term. The second term is a regularisation expression (whose strength is controlled by the hyperparameter $\rho$) penalising the mean squared deviation between the adjusted turnover parameters and the original parameter set. The latter term mainly serves two purposes: first, it ensures that, whenever a turnover parameter is ”free” in the problem (which can happen if its associated reaction fluxes are 00 across all conditions, or if no experimental measurements are available for its associated enzyme), it will be kept at its original value; secondly, it provides a mean to tune the strength of the adjustment procedure.
In order to define the region $\mathcal{S}$, we shall exploit the sufficient optimality conditions of LP (28). Introducing, for each growth condition, the vectors of dual variables, $\boldsymbol{\lambda}_{j}\in\mathbb{R}^{M}$ and $\boldsymbol{\mu}_{j}\in\mathbb{R}^{P}$, corresponding to constraints 28b and 28c, respectively, we note that problem 28 admits (for the $j$th growth condition) a dual problem in the form:
@@
A well-known result in linear programming duality theory [maranas_zomorrodi_2016] states that a flux distribution is the optimal solution of the primal problem 28, if, and only if, the dual problem 34 is feasible (dual feasibility) and its optimal objective coincides with the primal optimal objective, that is $\mathbf{a}_{j}^{\top}\mathbf{v}_{j}^{*}=-\mathbf{b}_{j}^{\top}\boldsymbol{\mu}_{j}$ (strong duality). Taken together, dual feasibility and strong duality thus define the region of optimality in turnover parameter space of each reference flux distribution. Hence, we can integrate the above definition of $\mathcal{S}$ within problem 32 by introducing the two vectors of dual variables ($\boldsymbol{\lambda}_{j}$ and $\boldsymbol{\mu}_{j}$) for each condition as additional optimisation variables, and simultaneously enforcing the optimality for each reference distribution. By doing this, we obtain the final formulation of the nonlinear program for turnover number adjustment, which we solved to obtain the results shown in the main text: